OpenStudy (bloomlocke367):

Evaluate each infinite series that has a sum. (I only need to do one) it's in the comments.

OpenStudy (bloomlocke367):

\[\sum_{n=1}^{\infty} (-\frac{ 1 }{ 3 })^{n-1}\]

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

use \[\frac{1}{1-r}\]and you are done in one step

OpenStudy (misty1212):

you can do it in your head what is one plus one third?

OpenStudy (bloomlocke367):

4/3

OpenStudy (misty1212):

yes and the reciprocal of 4/3?

OpenStudy (bloomlocke367):

3/4

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (bloomlocke367):

Why'd you do that, though?

OpenStudy (bloomlocke367):

@misty1212

OpenStudy (misty1212):

what does the question "why" mean in this context? does it mean why is \[\sum_{n=1}^{\infty}r^{n-1}=\frac{1}{1-r}\]?

OpenStudy (misty1212):

you can derive the formula for yourself, and i am sure its derivation is in your book, it is not hard, but would be rather lengthy to write here

OpenStudy (bloomlocke367):

yes. Where did \[\frac{ 1 }{ 1-r }\] come from?

OpenStudy (bloomlocke367):

wait, nevermind... it comes from\[S=\frac{ a_1 }{ 1-r}\]

OpenStudy (misty1212):

that is how you sum a geometric series like this one outline of proof like this, if you call the sum \(S\) \[S=1+r+r^2+r^3+...\\ rS=r+r^2+r^3+r^4+...\\ S-rS=1\\ S(1-r)=1\\ S=\frac{1}{1-r}\]

OpenStudy (preetha):

Wow, very nicely done Misty! I am impressed!

OpenStudy (misty1212):

@Preetha thanks!

OpenStudy (bloomlocke367):

My book says that if \[\left| r \right|<1\] it has a finite sum and if\[\left| r \right|\ge1\] it doesn't have a finite sum.. so, since r was -1/3, there's a sum?

OpenStudy (bloomlocke367):

and that also means it converges, correct?

OpenStudy (misty1212):

oh yes that is because \[|-\frac{1}{3}|<1\]

OpenStudy (bloomlocke367):

Okay. Thank you so much for your clarification!